OpenStudy (anonymous):

A boy with a mass of 66 kg sits on a see-saw, 1.2 meters from the fulcrum. How far from the fulcrum will a 44-kg girl need to sit to balance the boy?

OpenStudy (abhisar):

\(\boxed{\boxed{\boxed{\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\color{blue}{\mathfrak{~~~~Welcome~to~OpenStudy!~\ddot\smile}}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}}}}\)

OpenStudy (abhisar):

Let the girl sit x metres from the fulcrum

OpenStudy (abhisar):

Then, 1.2*66*10 = 44*10*x, Solving this we get x=1.8 m

OpenStudy (abhisar):

got it ?

OpenStudy (anonymous):

haha yes thanks

OpenStudy (abhisar):

Did you know you can give medals to people who help you with questions? Just click the \(\color{blue}{\text{blue "best response" button}}\) to show appreciation for the people who help you.

OpenStudy (abhisar):

\(\Huge\text{Anytime !}\) \(\huge\ddot\smile\)